OpenStudy (anonymous):

A basketball team has won 50 games of 75 played. The team still has 45 games to play. How many of the remaining games must the team win in order to win 60% of all games played during the season?

OpenStudy (lilg132):

22 games more to win in order to win 60% of all games played during the season

OpenStudy (pokemon23):

what's up Lilg

OpenStudy (lilg132):

75 + 45 remaining games = 120 games in total 120 / 100 = 1.2 (this is 1 percent of the games) 1.2 * 60 = 72 games is 60 percent of all games remember to click good answer

OpenStudy (lilg132):

nothing much you

OpenStudy (pokemon23):

it ok :)

OpenStudy (anonymous):

Let x be the number games that have to be won beyond the 50 already won. solve the following for x:\[\frac{50+x}{75+45}=\frac{60}{100} \]\[x=22 \]